pls help me solve this question 6*(9/3)=2a , 15(3/b)=45/b , (c-3)+3=c ,
(a+b)-c=a+(b-c) , 7a+(4+5a)=12a+4 pls answer this thank u!

Answer by user_dude2008(1862)

6*(9/3)=2a
6*3=2a
18=2a
9=a
a=9



15(3/b)=45/b
45/b=45/b
b=any real number
Answer: infinite number of solutions


(c-3)+3=c
c+(-3+3)=c
c+0=c
c=c
c=any real number
Answer: infinite number of solutions


(a+b)-c=a+(b-c)
a+b-c=a+b-c
0=0
Answer: infinite number of solutions


7a+(4+5a)=12a+4
12a+4=12a+4
Answer: infinite number of solutions